a fire truck can drive a max of 40 miles and hour and is 4.5 miles away, how many minutes will it take it to get there?
Answer by stanbon(75887)

a fire truck can drive a max of 40 miles and hour and is 4.5 miles away, how many minutes will it take it to get there?
------
time = distance/rate
----
time = 4.5/40 = 0.1125 hrs = 0.1125(60 min) = 6.75 minutes
============================
Cheers,
Stan H.